A parallel plate capacitor is charged to 100v.its plate separation is 2mm and the area of separation of each plate is 120cm^2.calculate and account for the increase and decrease in energy when separation is reduced to 1mm at constant voltage and charge




1
Expert's answer
2022-01-21T08:25:05-0500

Given:

"V=100\:\rm V"

"d_1=0.002\:\rm m"

"d_2=0.001\:\rm m"

"A=0.012\:\rm m^2"

The initial and final capacities of capacitor

"C_1=\frac{\epsilon_0 A}{d_1}=\frac{8.85*10^{-12}*0.012}{0.002}=5.31*10^{-11}\:\rm F"

"C_2=\frac{\epsilon_0 A}{d_2}=\frac{8.85*10^{-12}*0.012}{0.001}=1.06*10^{-10}\:\rm F"

The change in energy of capacitor

"\Delta W=\frac{C_2-C_1}{2}V^2"

"=\frac{5.31*10^{-11}-1.06*10^{-10}}{2}100^2=-2.66*10^{-7}\rm\: J"